OpenStudy (ttp):

Can someone tell me if i'm correct?? A bicycle manufacturing company makes a particular type of bike. Each child bike requires 4 hours to build and 4 hours to test. Each adult bike requires 6 hours to build and 4 hours to test. With the number of workers, the company is able to have up to 120 hours of building time and 100 hours of testing time for a week. If c represents child bikes and a represents adult bikes, determine which system of inequality best explains whether the company can build 10 child bikes and 12 adult bikes in the week.

OpenStudy (ttp):

No, because the bike order does not meet the restrictions of 4c + 6a ≤ 120 and 4c + 4a ≤ 100 No, because the bike order does not meet the restrictions of 4c + 4a ≤ 120 and 6c + 4a ≤ 100 Yes, because the bike order meets the restrictions of 4c + 6a ≤ 120 and 4c + 4a ≤ 100 Yes, because the bike order meets the restrictions of 4c + 4a ≤ 120 and 6c + 4a ≤ 100

OpenStudy (ttp):

I think it's C.

OpenStudy (anonymous):

4c + 6a <= 120 [building hours] 4c + 4a <= 100 [testing hours] 4 * 20 + 6 * 6 <= 120 80 + 36 <= 120 116 <= 120 true, they have enough building hours 4 * 20 + 4 * 6 <= 100 80 + 24 <= 100 104 <= 100 false, they won't have enough testing hours the first option is the answer